OpenStudy (anonymous):

Find the solution to the equation 64^(4 - x) = 4^2x Using complete sentences, explain the procedure used to solve this equation

OpenStudy (nurali):

64^(4 - x) = 4^2x 4^3(4-x)=4^2x 3(4-x)=2x 12-3x=2x 12=5x x=12/5